# Poisson Hail on a Wireless Ground

### v1

@@ 5.2.2 Continuous-Space Setting
We consider exclusion sets $\{S_{i}\}$ of two types: (1) balls of random radius centered at the point of arrival, where the random radii are i.i.d. exponential random variables of mean $R$; (2) balls of fixed radius $R$ centered at the point of arrival. The heights $\{h_{i}\}$ are i.i.d. and exponentially distributed.
@@
We simulate the example defined on the torus version of ${\cal X}=[-W,W]^{2}$, $b=1,s=l(0)$. We compare the stability region (captured by the critical threshold) with different fixed or i.i.d. exponentially distributed radius for a fixed function $l$.
@@
We first give the analytical results of two special cases of exclusion sets. For $R\equiv 0$, the model is a WSBD process defined in [3]. The stability region is $\lambda>\lambda_{c}$, where the critical threshold is [3]
@@
On the other hand, the dynamic reduces to that of a M/M/1 queue with global FCFS when the exclusion sets are such that there is at most one customer on the ground at any time. For such an M/M/1 system, the stability region is also known as $\lambda>\lambda_{c}$, and
$\lambda_{c}=\frac{c}{|{\cal X}|\mathbb{E}h},$ (15)
@@
where $c$ is the constant denoting the maximum Shannon rate achieved when there is no interference as defined earlier.
@@
Figure 3: $\cal X$ = $[-2,2]^{2}$. $\mathbb{E}h=1$, i.i.d. exponential radius, and function $l(r)=\min(1,r^{-4}),w=0.05$.
Figure 4: $\cal X$ = $[-10,10]^{2}$. $\mathbb{E}h=1$, fixed constant radius, and function $l(r)=\min(1,r^{-4}),w=0.05$.
@@
In Fig. 3, we consider ${\cal X}=[-2,2]^{2}$, $\mathbb{E}h=1,w=0.05$, and function $l(r)=\min(1,r^{-4})$, with i.i.d. exponential radii. We plot the critical threshold versus the mean radius. From Eq. (14) and (15), the critical thresholds for the special cases $\mathbb{E}R=0$ and $\mathbb{E}R=\infty$ are 0.2558 and 0.2745 for $R=0$ and M/M/1, respectively. The estimation of the stability threshold is done as described in Subsection 5.1 with 1000,000 realizations of blocks (with the exception of the data point at $\mathbb{E}R=0.4$, which is calculated from 3,500 block realizations due to its system complexity).
In Fig. 4, we consider ${\cal X}=[-10,10]^{2}$, $\mathbb{E}h=1,w=0.05$, function $l(r)=\min(1,r^{-4})$, and fixed constant radius. Here, due to the complexity of the system, the stability threshold for the fixed radius at $R\equiv 0.5$ is estimated by comparing the delay distributions of the system with 100,000 arrivals under different arrival intensities.
In both cases, the critical threshold versus exclusion radius displays unimodality: $\lambda_{c}$ has a single peak and the largest value appears in the middle $0<R^{*}<\infty$. This leads to what we refer to as optimal exclusion. The results show that, for the given scenario, moderate protection optimizes the system stability as opposed to immediate access, which is too greedy, or global FIFO, which is too conservative.
6 Spatial Interaction Queuing Processes
@@
This section defines a new and general class of spatial queuing processes. Arrivals in these systems are in terms of e.g. a Poisson rain or a renewal rain as above. Arrivals are independently marked with a locus of arrival $x$ in a given compact $D$ of the Euclidean space, a height $h$ representing the size of the task, and an exclusion set $B$. Service is again locally FCFS, based on the exclusion sets, with the same definition as above. The generalization stems from the fact that the interaction function, which gives the service speed of customers in a given spatial configuration, is general. More precisely, let $\psi$ denote the point process of customers currently being served. For all $x\in\psi$, the service rate of $x$ in this configuration is of the form $s(x,\psi)$, where $s$ is a measurable function from $D\times{\cal M}_{D}$ to $\mathbb{R}^{+}$. Here, ${\cal M}_{D}$ denotes the space of counting measures on $D$. The Shannon rate function (1) used above was of this form. In this special case, the map is monotonic in $\psi$. In this section, we do not make any assumption of this kind. We assume that the $s$ function defined above is bounded below by a positive constant and bounded above by another constant.
@@
It can be seen that the induction of Section 3 extends without modifications to this more general setting. Similarly, blocks can be defined in the same way as in Subsection 4.3. Using this, one can show that the stability condition of this general class of systems is also of the form
@@
(with the notation of Section 4), with the same interpretation of $\widehat{\sigma}$ as the time to serve a block when all elements of the said block are available at time zero and are served in the FCFS order prescribed by the constraints at the speed prescribed by the interaction function $s$.
7 Discussion and Open Questions
7.1 On the complexity of this class of systems
The first set of questions concerns the complexity of the estimation of $\lambda_{c}$ and of other statistical properties of this class of dynamical systems. If there are a few cases where one can get closed-form expressions for $\lambda_{c}$, even the simulation of a perfect sample of the two random variables allowing one to evaluate $\lambda_{c}$ (see Subsection 5.1) may be a challenging problem. This is well illustrated by the following discrete model: the loci of arrivals are uniformly distributed on the integers ranging from 1 to $N$ and the exclusion radii are equal to $2$ (namely a customer of locus $k$ can only be served after all customers of locus $k-2,k-1,k,k+1$ and $k+2$ that arrived before have left - conditions involving integers not in $[1,N]$ are void). Assuming that $N$ is even, a natural collection of sets implementing the conditions of Subsection 4.1 is $B_{1}$ the ball (of diameter 3) centered in 2, $B_{2}$ that centered in 4, etc, up to $B_{K}$ with $K=\frac{N-1}{2}$ that centered in $N-1$. Then the probability of event $C$ is
$\left(\frac{1}{N}\right)^{K}.$
For $N=21$, the mean number of arrivals until event $C$ is $21^{10}>10^{13}$.
7.2 Connection to the Saturation Rule
The results obtained in Section 6 can be seen as a partial extension of the saturation rule of [11] to nonmonotonic dynamics (in contrast to PHHG for instance). For all models of spatial interaction queues considered there, the stability region is obtained when assessing the time to empty the system when blocking arrivals, exactly as in the saturation rule.
7.3 Open problems
@@
    The saturation rule was established in the stationary ergodic framework. In contrast, the results of the present paper require independent assumptions. Can one relax some of these assumptions and still be able to build the stationary regime?
@@
    The Poisson Hail model was successfully defined on the infinite Euclidean space. Can one construct the PHWG model and extend the stability results obtained in the present paper to non-compact phase spaces? On most sensible models, there will be no zigzag event of positive probability. Therefore, another mathematical approach will be needed.
@@
    The basic dynamics is fundamentally nonmonotonic in that delaying the arrival of a given customer may lead to an earlier departure for another customer. This is why the saturation rule cannot be applied. Can one circumvent this difficulty, e.g., by averaging over permutations of the exclusion sets?
@@
    It would be practically useful to adapt the model described here to a situation closer to the CSMA/CA setting. This will in particular require the extension of this model to disciplines different from FCFS and more precisely to disciplines based on the carrier sensing and back-off mechanisms used in this context. Conversely, it would be interesting to devise concrete distributed algorithms that allow one to implement the model described in the present paper.
@@ Acknowledgments
8 Appendix
8.1 Discrete time
In this subsection, we reformulate our Theorem 1 in a simpler and more general form, as Proposition 13, and provide its proof.
Assume that, for each $x\geq 0$, there is given an i.i.d. sequences $\{\xi_{n}(x)\}$ of random variables with finite mean ${\mathbb{E}}\xi_{1}(x)$, and that all these sequences are mutually independent. Introduce the following recurrence relation: $X_{0}=const$ a.s. and
$X_{n+1}=(X_{n}+\xi_{n}(X_{n}))^{+},\ \ n\geq 1.$ (16)
The sequence $\{X_{n}\}$ is a time-homogeneous Markov chain, and the following result holds.
@@ 8.1 Discrete time
Assume that the family $\{\xi_{1}(x)\}_{x\geq 0}$ is uniformly integrable and that
$\xi_{1}(x)\rightarrow_{st}\xi\equiv\xi(\infty),\ \text{ as }\ x\to\infty.$ (17)
@@
    If ${\mathbb{E}}\xi<0$, then there exists a finite $K>0$ such that the set $[0,K]$ is positive recurrent, i.e., $\tau_{x}\equiv\tau_{x}(K)=\min\{n\geq 1:X_{n}\leq K\}<\infty$ a.s., for any $X_{0}=x\geq 0$ and, moreover,
@@
    If, in addition, the Markov chain has a positive recurrent atom at point, say, $x_{0}$ and if ${\mathbb{P}}(X_{1}=x_{0}\ |\ X_{0}=x_{0})>0$, then there exist a unique stationary distribution, say $\pi$, and
@@
    in the total variation, for any initial value $X_{0}=x$.
@@
    If ${\mathbb{E}}\xi_{1}>0$ and if there is a constant $\gamma>0$ such that, for any $x$,
@@
[…]
@@
In the case where ${\mathbb{E}}\xi=0$, there may be various scenarios, including positive recurrence, null recurrence and transience – depending on the way the distribution of $\xi_{1}(x)$ converges to the distribution of $\xi$ as $x$ grows.
@@
Consider a time-homogeneous Markov chain $\{X_{n}^{(x)}\}$ on a measurable state space $({\cal X,B_{X}})$ that starts from initial state $X_{0}^{(x)}=x$. Let $L:{\cal X}\to\infty$ be an unbounded measurable function. Let
@@
Assume there exists $\varepsilon>0$ and $K>0$ such that
@@
We let $\xi_{n}(x)=\sigma_{n}(x)-t_{n}$. Further, let $\varepsilon:=-{\mathbb{E}}\xi>0$ and $\Delta(x)={\mathbb{E}}(X_{1}-X_{0}\ |\ X_{0}=x)$. We have
@@
Further, the existence of a positive recurrent atom and aperiodicity imply the stability result.
In the particular case of Theorem 1, the two latter properties are guaranteed by unboundedness of the support of the distribution of interarrival times.
Before giving the proof of the statement (ii) of Proposition 13, we recall Theorem 35 in Section 11 of [12] which is formulated below. Its proof is a minor modification of the proof of the corresponding result from PhD thesis by [13]. and, in turn, the latter result is a slight generalization of the instability theorem in [14].
Consider a time-homogeneous Markov chain $\{X_{n}^{(x)}\}$ on a measurable state space $({\cal X,B_{X}})$ that starts from initial state $X_{0}^{(x)}=x$. Let $L:{\cal X}\to\infty$ be an unbounded measurable function. For $K>0$, let
$\tau_{x}(K)=\min\{n\geq 1\ :\ X_{n}^{(x)}>K\}.$
@@
[…]
@@
Here $({\cal X,B_{X}})$ is the positive half-line with the Borel sigma-algebra and $L(x)=x$. We again let $\xi_{n}(x)=\sigma_{n}(x)-t_{n}$. Then condition 1 of the theorem follows from condition (13), condition (21) from the uniform integrability of $\{\xi_{1}(x),x\geq 0\}$, and condition (17), and condition (22) follow from the fact that the random variables $\{t_{n}\}$ are i.i.d. with a finite mean, thus one can take $\eta=t_{1}$. ∎
8.2 Continuous time, proof of Theorem 2
Proof of (i). This follows from known results for regenerative processes – see e.g. Theorem 6.1.2 and Proposition 7.3.8 in [15].

### v2

@@ IV-B1 Continuous-Space Setting
We consider exclusion sets $\{S_{i}\}$ of two types: (1) balls of random radii centered at the point of arrival, where the random radii are i.i.d. exponential random variables of mean $R$; (2) balls of fixed radius $R$ centered at the point of arrival. In both cases, the heights $\{h_{i}\}$ are i.i.d. and exponentially distributed.
@@
We simulate the example defined on the torus version of ${\cal X}=[-Q,Q]^{2}$, $b=1,s=l(0)$. We compare the stability region (captured by the critical threshold) with different fixed or i.i.d., exponentially distributed radii, for a fixed function $l$.
@@
We first give the analytical results on the two special cases of exclusion sets. For $R\equiv 0$, the model is a WSBD process defined in [17]. The stability region is $\lambda<\lambda_{c}$, where the critical threshold is
@@
On the other hand, the dynamic reduces to that of an M/M/1 queue with global FCFS when the exclusion sets are such that there is at most one customer on the ground at any time. For such an M/M/1 system, the stability region is also known as $\lambda<\lambda_{c}$, and
$\lambda_{c}=\frac{c}{\mathbb{E}h},$ (16)
@@
where $c$ is the constant denoting the maximum Shannon rate achieved when there is no interference, as defined earlier.
@@
Figure 4: $\cal X$ = $2,2]^{2}$. $\mathbb{E}h=1$. The exclusion sets have i.i.d. exponential radius, and function $l(r)=\min(1,r^{-4}),w=0.05$.
Figure 5: The log ccdf of delay (time elapsed between the arrival and departure of customers) for different constant radii. $\cal X$ = $10,10]^{2}$. $\lambda=0.15$. $\mathbb{E}h=1$ and $l(r)=\min(1,r^{-4}),w=0.05$.
@@
In Fig. 4, we consider ${\cal X}=[-2,2]^{2}$, $\mathbb{E}h=1,w=0.05$, and the function $l(r)=\min(1,r^{-4})$, with i.i.d. and exponential radii for the exclusion sets. We plot the critical threshold versus the mean radius. From Eq. (15) and (16), the critical thresholds for the special cases $\mathbb{E}R=0$ and $\mathbb{E}R=\infty$ are 0.2558 and 0.2745 for $R=0$ and M/M/1, respectively. The estimation of the stability threshold is done as described in Subsection IV-A with 1000,000 realizations of blocks (with the exception of the data point at $\mathbb{E}R=0.4$, which is calculated from 5,500 block realizations due to the complexity of the system. The stability threshold, seen as a function of the mean exclusion radius, displays unimodality: $\lambda_{c}$ has a single maximum which appears at $0<E[R]<\infty$. This leads to what can be referred to as an “optimal exclusion radius”. The results show that, for the given scenario, moderate protection optimizes the system stability as opposed to immediate access, which is too greedy, or global FIFO, which is too conservative.
In Fig. 5, we compare the delay distributions of the system with 1000,000 arrivals and arrival intensity $\lambda=0.15$ for different constant radii, $R=0,0.5,0.8,1$. ${\cal X}=[-10,10]^{2}$, $\mathbb{E}h=1,w=0.05$, $l(r)=\min(1,r^{-4})$. The case with $R=0$ is equivalent to WSBD. From Eq. (15) and (16), the critical thresholds for the special cases $R=0$ and $R\geq 5.8579$ are 0.23 and 0.011 for $R=0$ and M/M/1, respectively. For $R=0.5$, simulations suggest that $\lambda_{c}\approx 0.25$. The delay distribution first improves from $R=0$ to $R=0.8$, and then decreases at $R=1$. For constant radii, the delay distributions are consistent with the unimodality of the stability threshold shown in Fig. 4. It is also worth noting that the delay distributions for $R=0$ and $R=1$ have different slopes that intersect each other at one point. This is likely because for $R=1$, the service rate is (significantly) higher than $R=0$ on average due to protection while the waiting time is (significantly) longer.
IV-B2 Discrete-Space Setting
We now define and analyze an example in the 1D discrete-space setting. Let $N$ be an even natural number $N\geq 4$. This model belongs to the general framework of the paper, with discrete state space and Poisson arrivals. There are $N$ arrival loci located equidistantly on the circle of length $N$. We number them from $1$ to $N$ clockwise, say. Customers are represented as (open) rectangles $M\times h$ centered at $x$. Here, $x$ is one of the loci. Here, $M$ is a fixed integer, with $1\leq M\leq N$. As in the previous discussions, we assume here that customers are served in a locally FCFS way. Finally, $h$ is a positive random variable which represents the file size of a customer.
The distance between two customer centers is the minimum of the two distances between them, clockwise and counterclockwise. Let $d$ be a fixed positive integer. The attenuation function at distance $u\in\mathbb{N}$ is $l(u)=1_{u=0}+v\cdot 1_{0<u\leq d}$, where $v>0$ is a constant. The thermal noise is equal to 1. Hence, the service rate of a customer at $x$ and in service is equal to ${1}/{(1+v\phi_{d}(x)})$ with $\phi_{d}(x)$ the number of customers being served at a distance less than or equal to $d$, without counting the customer at $x$.
Note that for $M>d$, this model coincides with Poisson Hail since the rate of any customer being served is always 1. It is known that Poisson Hail is strictly monotonic in the footprint of the hailstones. More precisely, for $N$ and $d$ fixed, the value of $\lambda_{c}$ is a decreasing function of $M$, in the range $M>d$.
Note that if $M>N/2$, then there is no way to serve two or more customers simultaneously, and if $N/2\geq M>N/3$, then there may be the simultaneous service of at most two customers.
A special case of this model is obtained by taking $d=N/2$ and $N/2\geq M>N/3$. That is, a customer’s service “interferes” with at least another customer, then it interferes with exactly one customer, and its service rate is then 1/(v+1){1}/{(v+1)}. If $M>N/2$, then there is no way to serve two or more customers simultaneously, and we have a single-server queue with service rate 1. If $M=N/2$, then the simultaneous service of two customers may only occur if the distance between the customer centers is exactly $N/2$, and at most two customers are served simultaneously, with rate 1/(v+1).{1}/{(v+1)}.
Consider a particular case $M=d=N/2$. Consider the saturated system, namely the system with infinitely many customers in the queue in front of it, which are numbered $1,2,\ldots$. Assume that customer $i$ is centered at $x_{i}$ and brings $h_{i}$ units of work to serve, that $\{x_{i}\}$ is an i.i.d. sequence with elements uniformly distributed on the discrete set of $N$ locations, and that $\{h_{i}\}$ is an i.i.d. sequence of positive random variables. Let us number customers $0,1,2,\ldots$ Assume that customer 0 has its center at $x_{0}$, and let $\nu_{1}$ be the number of the first customer whose center is located neither at $x_{0}$ nor at $(x_{0}+N/2)\mbox{mod}\ N$. Then we define the first block of customers as customers enumerated $0,1,\ldots,\nu_{1}-1$, and their number, $\nu_{1}$, has a geometric distribution with parameter (N2)/N{(N-2)}/{N}. Further, each customer in the block is centered at $x_{0}$ with probability $1/2$, independently of anything else. All following blocks are formed similarly. So customers are served in blocks of random sizes that are i.i.d. and have a geometric distribution with parameter p=(N2)/Np=(N-2)/N. Given that there are, say $\nu=n$ customers in a “typical” block, a binomial number of them, say $m_{n}\sim Bin(n,1/2)$, have the total amount of work $H_{1}:=\sum_{1}^{m_{n}}h_{i}$ and occupy the same $N/2$ servers and $n-m_{n}$ others occupy the other $N/2$ servers and have the total amount of work $H_{2}:=\sum_{1}^{n-m_{n}}h^{\prime}_{j}$ where $\{h^{\prime}_{j}\}$ are independent copies of $h_{1}$. Then the mean service time of a typical block of customers, say $s$, is
$s$ $={(1+v)}{\mathbb{E}}\min(H_{1},H_{2})+{\mathbb{E}}|H_{2}-H_{1}|$
$\stackrel{{\scriptstyle\text{\scriptsize(a)}}}{{=}}{(1+v)}{\mathbb{E}}\min(H_{1},H_{2})+{\mathbb{E}}(H_{1}+H_{2}-2\min(H_{1},H_{2}))$
$\stackrel{{\scriptstyle\text{\scriptsize(b)}}}{{=}}\frac{N}{N-2}{\mathbb{E}}h_{1}+{(v-1)}{\mathbb{E}}\min(H_{1},H_{2}),$
where step (a) follows from the fact that $|H_{2}-H_{1}|=\max(H_{1},H_{2})-\min(H_{1},H_{2})=H_{1}+H_{2}-2\min(H_{1},H_{2})$, and step (b) follows from the linearity of expectation and the mean of the geometric distribution of the number of customers with parameter (N2)/N(N-2)/N.
In the particular case where $h_{i}=1$ a.s., the formula for $s$ becomes slightly simpler:
$s=\frac{N}{N-2}+(v-1){\mathbb{E}}\min(m_{\nu},\nu-m_{\nu}).$
Here the critical value $\Lambda_{c}$ is
$\Lambda_{c}=\frac{N}{(N-2)s.}$
V Spatial Interaction Queuing Processes
@@
This section defines a new and general class of spatial queuing processes. Arrivals in these systems are in terms of e.g. a Poisson rain or a renewal rain, as above. Arrivals are independently marked, with a locus of arrival $x$ in a given compact $D$ of the Euclidean space, a height $h$ representing the size of the task, and an exclusion set $B$. The service discipline is locally FCFS, based on the exclusion sets, with the same definition as above. The generalization comes from the fact that the interaction function, which gives the service speed of customers in a given spatial configuration, is general. More precisely, let $\psi$ denote the point process of customers currently being served. For all $x\in\psi$, the service rate of $x$ in this configuration is of the form $s(x,\psi)$, where $s$ is a measurable function from $D\times{\cal M}_{D}$ to $\mathbb{R}^{+}$. Here, ${\cal M}_{D}$ denotes the space of counting measures on $D$. The Shannon rate function (2) used above was of this form. In this special case, the map is monotonic in $\psi$. In this section, we do not make any assumption of this kind. We assume that the $s$ function defined above is bounded below by a positive constant and bounded above by another constant.
@@
It can be seen that the induction of Section II extends without modifications to this more general setting. Similarly, blocks can be defined in the same way as in Subsection III-C. Using this, one can show that the stability condition of this general class of systems is also of the form
@@
(with the notation of Section III), with the same interpretation of $\widehat{\sigma}$ as the time to serve a block when all elements of the said block are available at time zero and are served in the FCFS order prescribed by the constraints at the speed prescribed by the interaction function $s$.
VI Discussion and Open Questions
VI-A On the Complexity of this Class of Systems
The first set of questions concerns the complexity of the estimation of $\lambda_{c}$ and of other statistical properties of this class of dynamical systems. If there are a few cases where one can get closed-form expressions for $\lambda_{c}$, even the simulation of a perfect sample of the two random variables allowing one to evaluate $\lambda_{c}$ (see Subsection IV-A) may be a challenging problem. This is well illustrated by the following discrete model: the arrival locations are uniformly distributed over the integers ranging from 1 to $N$ on the real line, and the exclusion radii are equal to $2$ (namely a customer of locus $k$ can only be served after all customers of locus k2,k1,k,k+1k-2,k-1,k,k+1 and $k+2$ that arrived before have left - conditions involving integers not in $[1,N]$ are void). Assuming that $N$ is odd, a natural collection of sets implementing the conditions of Subsection III-A1 is $B_{1}$ the ball (of diameter 3) centered in 2, $B_{2}$ that centered in 4, etc, up to $B_{K}$ with $K=\frac{N-1}{2}$ that centered in N1N-1. Then the probability of event $C$ is $\left(\frac{1}{N}\right)^{2K}.$ For $N=21$, the mean number of arrivals until event $C$ is $21^{20}>10^{26}$.
VI-B Connection to the Saturation Rule
The results obtained in Section V can be seen as a partial extension of the saturation rule of [3] to non-monotonic dynamics (in contrast to PHHG for instance). For all models of spatial interaction queues considered there, the stability region is obtained when assessing the time to empty the system when blocking arrivals, exactly as in the saturation rule.
VI-C Open Problems
@@
    The saturation rule was established in the stationary ergodic framework. In contrast, the results of the present paper require independence assumptions. Can one relax some of these assumptions and still be able to build the stationary regime?
@@
    The Poisson Hail model was successfully defined on the infinite Euclidean space. Can one construct the PHWG model and extend the stability results obtained in the present paper to non-compact phase spaces? On such spaces, for most sensible models, there will be no zigzag event of positive probability. Therefore, another mathematical approach will be needed.
@@
    The basic dynamics is fundamentally non-monotonic in that delaying the arrival of a given customer may lead to an earlier departure for another customer. This is why the saturation rule cannot be applied. Can one circumvent this difficulty, e.g., by averaging over permutations of the exclusion sets?
@@
    It would be practically useful to adapt the model described here to a situation closer to the CSMA/CA setting. This will in particular require the extension of this model to disciplines different from FCFS, and more precisely to disciplines based on the carrier sensing and back-off mechanisms used in this context. Conversely, it would be interesting to devise concrete distributed algorithms that allow one to implement the model described in the present paper.
@@ Acknowledgments
VII Appendix
VII-A Discrete time
For $x\geq 0$ and $n=1,2,\ldots$, let
$\xi_{n}(x)=\Sigma_{n}(x)-(T_{n+1}-T_{n}).$ (17)
It follows from the earlier observations that the families of random variables
$\{\xi_{1}(x),x\geq 0\},\{\xi_{2}(x),x\geq 0\},\ldots,\{\xi_{n}(x),x\geq 0\},\ldots\ \ \text{are i.i.d. in}\ \ n$ (18)
(where, for each $n$, the trajectory of $\xi_{n}(x)$ is piecewise linear, with a countable number of jumps). Then the recursion (10) may be rewritten in the form:
$X_{n+1}=(X_{n}+\xi_{n}(X_{n}))^{+},\ \ n\geq 1,$ (19)
with $X_{n}=W_{n}$, and it is a time-homogeneous Markov chain.
We reformulate our Theorem 1 in a simpler and more general form, as Proposition 15, and provide its proof. Then Theorem 1 will follow if we let $\xi_{n}(x)$ be defined as in (17).
@@ VII-A Discrete time
Let a Markov chain $\{X_{n}\}$ follow recursion (19), with condition (18) being in force. Assume that the family $\{\xi_{1}(x)\}_{x\geq 0}$ is uniformly integrable and that
$\xi_{1}(x)\rightarrow\xi\equiv\xi(\infty)\ \text{ in distribution, as }\ x\to\infty.$ (20)
@@
    Let ${\mathbb{E}}\xi<0$. Then there exists a finite $K>0$ such that the set $[0,K]$ is positive recurrent, i.e., $\tau_{x}\equiv\tau_{x}(K)=\min\{n\geq 1:X_{n}\leq K\}<\infty$ a.s., for any $X_{0}=x\geq 0$ and, moreover,
@@
    If, in addition, the Markov chain has a positive recurrent atom at point, say, $x_{0}$ and if ${\mathbb{P}}(X_{1}=x_{0}\ |\ X_{0}=x_{0})>0$, then there exists a unique stationary distribution, say $\pi$, and
@@
    in total variation, for any initial value $X_{0}=x$.
@@
    Let ${\mathbb{E}}\xi_{1}>0$. If there is a constant $\gamma>0$ such that, for any $x$,
@@
    and if there exists an integrable random variable, say $\eta>0$ a.s., such that, for any $x$,
    $\xi_{1}(x)\geq-\eta\quad\text{a.s.},$ (24)
@@
    $\liminf_{n\to\infty}\frac{X_{n}}{n}\geq\varepsilon\quad\text{a.s.}$
@@
In the case where ${\mathbb{E}}\xi=0$, there may be various scenarios, including positive recurrence, null-recurrence and transience – depending on the way the distribution of $\xi_{1}(x)$ converges to the distribution of $\xi$ as $x$ grows.
@@
Consider a time-homogeneous Markov chain $\{X_{n}^{(x)}\}$ on a measurable state space $({\cal X,B_{X}})$ that starts from the initial state $X_{0}^{(x)}=x$. Let $L:{\cal X}\to[0,\infty)$ be an unbounded measurable function. Let
@@
Assume that there exists $\varepsilon>0$ and $K>0$ such that
@@
Let $\varepsilon:=-{\mathbb{E}}\xi>0$ and $\Delta(x)={\mathbb{E}}(X_{1}-X_{0}\ |\ X_{0}=x)$. We have
@@
Furthermore, the existence of a positive recurrent atom and aperiodicity imply the stability result. ∎
Remark 20.
In the particular case of Theorem 1, the existence of a positive recurrent atom and aperiodicity are guaranteed by the unboundedness of the support of the exponential distribution of interarrival times.
Before giving the proof of statement (ii) of Proposition 15, we recall Theorem 35 in Section 11 of [10] which is formulated below. Its proof is a minor modification of the proof of the corresponding result from PhD thesis by [8]. and, in turn, the latter result is a slight generalization of the instability theorem in [11].
Consider a time-homogeneous Markov chain $\{X_{n}^{(x)}\}$ in a measurable state space $({\cal X,B_{X}})$ that starts from the initial state $X_{0}^{(x)}=x$. Let $L:{\cal X}\to[0,\infty)$ be an unbounded measurable function. For $K>0$, let
$\theta_{x}(K)=\min\{n\geq 1\colon L\left(X_{n}^{(x)}\right)>K\}.$
@@
$\liminf_{n\to\infty}\frac{L(X_{n}^{(x)})}{n}\geq\varepsilon\quad\text{a.s.}$
@@
Here $({\cal X,B_{X}})$ is the positive half-line with the Borel sigma-algebra and $L(x)=x$. Then condition 1 of Theorem 3 follows from condition (23), condition (25) from the uniform integrability of $\{\xi_{1}(x),x\geq 0\}$ and from condition (20), and condition (26) is condition (24). ∎
Remark 21.
In the particular case of Theorem 1, condition (23) is condition (14) and condition (24) follows from the fact that the random variables $\{T_{n+1}-T_{n}\}$ are i.i.d. with a finite mean, thus one can take $\eta=T_{2}-T_{1}$.
VII-B Continuous time, proof of Theorem 2
Proof of (i). This follows from known results for regenerative processes – see e.g. Theorem 6.1.2 and Proposition 7.3.8 in [1] or Theorem 19.1 in Chapter 4 or [5] or Theorem 1 in Chapter 7 of [4].